OpenStudy (anonymous):

differentiate the following: y=w^3/2(w+5we^w)

OpenStudy (anonymous):

use quotient rule

OpenStudy (anonymous):

Yeah I know its just that I keep getting stuck

OpenStudy (anonymous):

it should be like \[\Large y'=\frac{1}{2}(\frac{\frac{d}{dw}(w^3)(w+5we^w)-\frac{d}{dw}(w+5we^w)*w^3}{(w+5we^w)^2})\]

OpenStudy (anonymous):

can you do this now ?

OpenStudy (anonymous):

Yeah I think the 5we^w was what was confusing me.Thanks a bunch :)

OpenStudy (anonymous):

where \[\Large \frac{d}{dw}(w^3)=3w^2\] and \[\Large \frac{d}{dw}(w+we^w)=(1+e^w+we^w)\]

OpenStudy (anonymous):

you are welcome :) and there is 5 i missed at the end !

OpenStudy (anonymous):

oh ok :)